# PEP-GS: Perceptually-Enhanced Precise Structured 3D Gaussians for View-Adaptive Rendering

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has achieved significant success in real-time, high-quality 3D scene rendering. However, it faces several challenges, including Gaussian redundancy, limited ability to capture view-dependent effects, and difficulties in handling complex lighting and specular reflections. Additionally, methods that use spherical harmonics for color representation often struggle to effectively capture specular highlights and anisotropic components, especially when modeling view-dependent colors under complex lighting conditions, leading to insufficient contrast and unnatural color saturation. To address these limitations, we introduce PEP-GS, a perceptually-enhanced framework that dynamically predicts Gaussian attributes, including opacity, color, and covariance. We replace traditional spherical harmonics with a Hierarchical Granular-Structural Attention mechanism, which enables more accurate modeling of complex view-dependent color effects and specular highlights. By employing a stable and interpretable framework for opacity and covariance estimation, PEP-GS avoids the removal of essential Gaussians prematurely, ensuring a more accurate scene representation. Furthermore, perceptual optimization is applied to the final rendered images, enhancing perceptual consistency across different views and ensuring high-quality renderings with improved texture fidelity and fine-scale detail preservation. Experimental results demonstrate that PEP-GS outperforms state-of-the-art methods, particularly in challenging scenarios involving view-dependent effects, specular reflections, and fine-scale details.
@@ 1 Introduction
Although 3D-GS demonstrates impressive rendering performance, it struggles with capturing complex view-dependent effects, especially in dynamic lighting environments. This limitation stems from the inherent constraints of low-order spherical harmonics (SH) used for color representation in 3D-GS, which are well-suited for modeling subtle, view-dependent phenomena but struggle to accurately represent specular highlights and anisotropic components. Due to their low-frequency nature, SH functions struggle to capture high-frequency, view-dependent details under dynamic lighting, which limits their effectiveness in complex scenes. Moreover, structured approaches like Scaffold-GS reduce Gaussian redundancy through hierarchical and region-aware scene representation but struggle with fine-scale details and texture modeling, leading to perceptual inconsistencies under complex lighting and reflections.
To address these challenges, we propose PEP-GS, a perceptually-enhanced framework for structured 3D Gaussian splatting inspired by previous work Lu et al. (2024); Laparra et al. (2016); Xu and Wan (2024). Similarly, we construct a sparse grid of anchor points, initialized from SfM points. For each neural Gaussian tethered to these anchors, PEP-GS dynamically predicts Gaussian attributes including opacity for transparency control, color for view-dependent appearance modeling, and covariance for spatial representation using a stable and precise framework. Specifically, we replace the traditional SH-based color representation with a Hierarchical Granular-Structural Attention mechanism, enabling more accurate modeling of view-dependent color effects and specular highlights. Additionally, we integrate Kolmogorov-Arnold Networks (KAN) for opacity and covariance estimation, enhancing the stability of the predictions. Building on this, we apply perceptual optimization techniques to improve the consistency of the rendered results. This approach refines the image details and texture fidelity, optimizing visual consistency across different views and significantly improving perceptual fidelity and fine-scale detail preservation in the final rendered images.
Through extensive experimentation, we demonstrate that our approach significantly outperforms existing state-of-the-art methods. As shown in Figure. 1, it shows particular strength in handling challenging scenarios such as specular reflections, view-dependent lighting effects, and intricate geometric details. Notably, our method maintains stability and consistently improves visual quality and perceptual consistency across novel views.
@@
  • Comprehensive experimental validation demonstrating substantial improvements in rendering quality across multiple datasets, particularly in challenging scenarios involving complex lighting, specular reflections, and fine geometric details.
@@ 3.2.1 View-Adaptive Color Refinement with Attention Mechanisms
To capture complex view-dependent effects such as specular reflections and highlights, we propose Hierarchical Granular-Structural Attention (HGSA), which integrates fine-grained feature enhancement and structural-level attention to dynamically aggregate features. This hierarchical mechanism adapts to directional changes and effectively combines detail-sensitive and global context information to optimize color representation.
Given an input feature $\mathbf{f}_{in}\in\mathbb{R}^{N\times(32+3+1)}$, as shown in Eq. 5, we first apply average pooling separately along two directions: horizontally (along the height) and vertically (along the width). These pooled features are then processed through a 1D convolutional layer, followed by a custom normalization and a sigmoid activation function to generate positional attention maps in both directions $y^{h}$ and $y^{w}$.
@@ 4.2 Comparisons
Table. 1 presents the results on three real-world datasets Barron et al. (2022); Knapitsch et al. (2017); Hedman et al. (2018), highlighting that our approach achieves competitive performance in rendering quality compared to state-of-the-art methods and stands out in terms of image quality and perceived similarity. While Scaffold-GS achieves the highest PSNR performance on the Deep Blending dataset, our method demonstrates more consistent and robust performance across other datasets, with PEP-GS notably outperforming Scaffold-GS on the Tanks&Temples dataset, as illustrated in Figure. 3. It can be noticed that our approach achieves comparable results with 3D-GS on Mip-NeRF360 dataset, and achieves superior results in Tanks&Templates datasets, exhibiting notable competitiveness in terms of SSIM and LPIPS, which captures more challenging environments with the presence of e.g.specular reflections, saturation and brightness under the bright light and thin structures.
@@
Figure 4: Compared to existing baselines, Our method demonstrates superior detail preservation, reduced artifacts, improved color consistency, and better handling of specular reflections across various scenes.
@@
Qualitative results across various datasets are provided in Figure. 4. We also conducted a more comprehensive and detailed comparison of rendering results against the scaffold method. As shown in Figure. 6, it is evident that our method demonstrates sharper details and textures, effectively preserving local structures and fine-scale details during image reconstruction. As shown in Figure. 7, under challenging sunlight environments, the Scaffold-GS method exhibits slightly excessive brightness, leading to the loss of certain details. In contrast, our method not only accurately maintains clearer textures and structural features but also avoids overexposure or shadow detail loss. Figure. 1 further illustrate that in challenging scenes like grass fields, our method better suppresses artifacts and unnatural textures while adeptly capturing complex textures and thin structures, exhibiting greater robustness. In terms of metric analysis, our method achieves lower LPIPS values and higher SSIM scores, confirming its significant advantage in overall visual quality and perceptual consistency.
@@ 4.3 Ablation Study
Figure 6: Comparison of fine-scale details in the DrJohnson scene within the Deep Blending dataset.
Figure 7: Comparison of brightness and saturation in the train scene within the Tanks&Templates dataset.
5 Conclusion
In this paper, we propose PEP-GS, a perceptually-enhanced framework for structured 3D Gaussian splatting, addressing key challenges in view-adaptive rendering. Our method introduces a Hierarchical Granular-Structural Attention mechanism to better capture complex view-dependent effects, and utilizes Kolmogorov-Arnold Networks for robust and interpretable Gaussian attribute estimation. Additionally, the incorporation of the Normalized Laplacian Pyramid Distance loss enhances perceptual consistency across views, ensuring high-quality renderings with superior texture fidelity and geometric precision. Experimental results show that our approach surpasses existing methods, excelling in handling complex lighting and fine-scale details, while maintaining rendering stability and efficiency.
@@ 5 Conclusion
We present comprehensive evaluation metrics for all scenes from the three public datasets mentioned in the main text. As shown in Tables. 3 and 4, our model consistently achieves superior performance across the majority of scenes. These quantitative results align with the enhanced perceptual capabilities of our model, further illustrated in Figures. 8 and 9.
@@
Figure 8: Comparison of texture-less regions and thin structures in the playroom scene within the Deep Blending dataset.

### v3

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has achieved significant success in real-time, high-quality 3D scene rendering. However, it faces several challenges, including Gaussian redundancy, limited ability to capture view-dependent effects, and difficulties in handling complex lighting and specular reflections. Additionally, methods that use spherical harmonics for color representation often struggle to effectively capture anisotropic components, especially when modeling view-dependent colors under complex lighting conditions, leading to insufficient contrast and unnatural color saturation. To address these limitations, we introduce PEP-GS, a perceptually-enhanced framework that dynamically predicts Gaussian attributes, including opacity, color, and covariance. We replace traditional spherical harmonics with a Hierarchical Granular-Structural Attention mechanism, which enables more accurate modeling of complex view-dependent color effects. By employing a stable and interpretable framework for opacity and covariance estimation, PEP-GS avoids the removal of essential Gaussians prematurely, ensuring a more accurate scene representation. Furthermore, perceptual optimization is applied to the final rendered images, enhancing perceptual consistency across different views and ensuring high-quality renderings with improved texture fidelity and fine-scale detail preservation. Experimental results demonstrate that PEP-GS outperforms state-of-the-art methods, particularly in challenging scenarios involving view-dependent effects and fine-scale details.
@@ 1 Introduction
Although 3D-GS demonstrates impressive rendering performance, it struggles with capturing complex view-dependent effects, especially in dynamic lighting environments. This limitation stems from the inherent constraints of low-order spherical harmonics (SH) used for color representation in 3D-GS, which are well-suited for modeling subtle, view-dependent phenomena but struggle to accurately represent anisotropic components. Due to their low-frequency nature, SH functions struggle to capture high-frequency, view-dependent details under dynamic lighting, which limits their effectiveness in complex scenes. Moreover, structured approaches like Scaffold-GS reduce Gaussian redundancy through hierarchical and region-aware scene representation but struggle with fine-scale details and texture modeling, leading to perceptual inconsistencies under complex lighting and reflections.
To address these challenges, we propose PEP-GS, a perceptually-enhanced framework for structured 3D Gaussian splatting inspired by previous work Lu et al. (2024); Laparra et al. (2016); Xu and Wan (2024). Similarly, we construct a sparse grid of anchor points, initialized from SfM points. For each neural Gaussian tethered to these anchors, PEP-GS dynamically predicts Gaussian attributes including opacity for transparency control, color for view-dependent appearance modeling, and covariance for spatial representation using a stable and precise framework. Specifically, we replace the traditional SH-based color representation with a Hierarchical Granular-Structural Attention mechanism, enabling more accurate modeling of view-dependent color effects. Additionally, we integrate Kolmogorov-Arnold Networks (KAN) for opacity and covariance estimation, enhancing the stability of the predictions. Building on this, we apply perceptual optimization techniques to improve the consistency of the rendered results. This approach refines the image details and texture fidelity, optimizing visual consistency across different views and significantly improving perceptual fidelity and fine-scale detail preservation in the final rendered images.
Through extensive experimentation, we demonstrate that our approach significantly outperforms existing state-of-the-art methods. As shown in Figure. 1, it shows particular strength in handling challenging scenarios such as view-dependent lighting effects, and intricate geometric details. Notably, our method maintains stability and consistently improves visual quality and perceptual consistency across novel views.
@@
  • Comprehensive experimental validation demonstrating substantial improvements in rendering quality across multiple datasets, particularly in challenging scenarios involving complex lighting and fine geometric details.
@@ 3.2.1 View-Adaptive Color Refinement with Attention Mechanisms
To capture complex view-dependent appearance changes under varying lighting and viewing conditions, we propose Hierarchical Granular-Structural Attention (HGSA), which integrates fine-grained feature enhancement and structural-level attention to dynamically aggregate features. This hierarchical mechanism adapts to directional changes and effectively combines detail-sensitive and global context information to optimize color representation.
Given an input feature $\mathbf{f}_{in}\in\mathbb{R}^{N\times(32+3+1)}$, which is formed by concatenating the viewing direction $\delta_{vc}$, the relative distance $\vec{\mathbf{d}}_{vc}$, and the integrated anchor feature $\hat{f}_{v}$:
$\mathbf{f}_{in}=\delta_{vc}\oplus\vec{\mathbf{d}}_{vc}\oplus\hat{f}_{v},$ (5)
where $\oplus$ denotes the operation of concatenation. As shown in Eq. 6, we first apply average pooling separately along two directions: horizontally (along the height) and vertically (along the width). These pooled features are then processed through a 1D convolutional layer, followed by a custom normalization and a sigmoid activation function to generate positional attention maps in both directions $y^{h}$ and $y^{w}$.
@@ 4.2 Comparisons
Table. 1 presents the results on three real-world datasets Barron et al. (2022); Knapitsch et al. (2017); Hedman et al. (2018), highlighting that our approach achieves competitive performance in rendering quality compared to state-of-the-art methods and stands out in terms of image quality and perceived similarity. While Scaffold-GS achieves the highest PSNR performance on the Deep Blending dataset, our method demonstrates more consistent and robust performance across other datasets, with PEP-GS notably outperforming Scaffold-GS on the Tanks&Temples dataset, as illustrated in Figure. 3. It can be noticed that our approach achieves comparable results with 3D-GS on Mip-NeRF360 dataset, and achieves superior results in Tanks&Templates datasets, exhibiting notable competitiveness in terms of SSIM and LPIPS.
@@
Figure 4: Compared to existing baselines, our method demonstrates superior detail preservation, reduced artifacts, and improved color consistency. Notably, in the second row of the images, our method also shows certain improvements in handling specular reflections compared to other approaches.
@@
Qualitative results across various datasets are provided in Figure. 4. We also conducted a more comprehensive and detailed comparison of rendering results against the scaffold method. As shown in the first and second rows of the left column in Figure. 6, it is evident that our method demonstrates sharper details and textures, effectively preserving local structures and fine-scale details during image reconstruction. As shown in the first and second rows of the right column in Figure. 6, under challenging sunlight environments, the Scaffold-GS method exhibits slightly excessive brightness, leading to the loss of certain details. In contrast, our method not only accurately maintains clearer textures and structural features but also avoids overexposure or shadow detail loss. Figure. 1 further illustrate that in challenging scenes like grass fields, our method better suppresses artifacts and unnatural textures while adeptly capturing complex textures and thin structures, exhibiting greater robustness. In terms of metric analysis, our method achieves lower LPIPS values and higher SSIM scores, confirming its significant advantage in overall visual quality and perceptual consistency.
@@
Figure 6: A systematic qualitative comparison between Scaffold-GS and PEP-GS is carried out across the Mip-NeRF360, Tanks&Temples, and Deep Blending datasets to highlight their relative strengths in diverse real-world scenarios.
@@ 4.3 Ablation Study
5 Conclusion and Limitation
In this paper, we propose PEP-GS, a perceptually-enhanced framework for structured 3D Gaussian splatting, addressing key challenges in view-adaptive rendering. Our method introduces a Hierarchical Granular-Structural Attention mechanism to better capture complex view-dependent effects, and utilizes Kolmogorov-Arnold Networks for robust and interpretable Gaussian attribute estimation. Additionally, the incorporation of the Normalized Laplacian Pyramid Distance loss enhances perceptual consistency across views, ensuring high-quality renderings with superior texture fidelity and geometric precision.
Even though experimental results show that PEP-GS surpasses existing methods, excelling in handling complex lighting and fine-scale details, there are still some limitations. Compared with scaffold-GS, while our method achieves superior metric scores and rendering performance, it slightly compromises rendering efficiency. We anticipate that our algorithm will progressively improve its rendering speed as the field advances.
6 Acknowledgments
This work is supported by the Hainan Provincial Natural Science Foundation of China (Grant No. 722RC678).
@@ 6 Acknowledgments
We present comprehensive evaluation metrics for all scenes from the three public datasets mentioned in the main text. As shown in Tables. 3 and 4, our model consistently achieves superior performance across the majority of scenes. These quantitative results align with the enhanced perceptual capabilities of our model, further illustrated in Figures. LABEL:fig:Comparison_of_texture-less_regions_and_thin_structures and 7.